envykaylie:

Jim earns a regular hourly rate of $5 and a premium hourly rate of $6. He worked 45 hours last week. What is his gross pay?

JeanLuc:

The premium is the overtime. Total time = 45 hours. 40 hours of regular time at $5, and 5 hours of overtime. Total = 40*5 + 5 * 6 = 200 + 30 = $230

JeanLuc:

Hope this helps!

JeanLuc:

Good luck with your quiz or whatever